OpenStudy (anonymous):

Write the equation of the parabola that is congruent to y=x^2: a) has a minimum at (-3,0) b) has a maximum at (0,7) c) has a minimum at (5,6)

OpenStudy (anonymous):

a ) y = x -3 b) y = x + 7 c) y = (x -5) + 6

OpenStudy (anonymous):

how did yu do it

OpenStudy (anonymous):

ok so you take the maximum/ minimum of your equation and that will be your value for y = (x -h) + k k is your vertical translation and h is your horizantal translation

OpenStudy (anonymous):

wudn't it be y=(x+3)+k

OpenStudy (anonymous):

oh my bad yeah the first one is y = (x -3) dont include the values if there are not given

OpenStudy (anonymous):

no its a pos 3 not a neg 3

OpenStudy (anonymous):

rem yu need to do it opp sign

OpenStudy (anonymous):

my bad again y = (x+ 3)

OpenStudy (anonymous):

LOL ok :D

OpenStudy (anonymous):

ty